A parallel plate condenser of 300uf is charged to 200volt . If the distance between its plates is halved what will be the potential difference between the plates and what will be the change in energy stored in it?

Grade:12

1 Answers

Adarsh
5 years ago
After charging , charge on each plates = 6 x 104micro columb.New capacitance when the distance between its plates is halved = 600uf. Potential difference between plates of capacitor = 60000/600 = 100v.Change in internal energy = ½.ci.vi2 – ½.cf.vf2 = 6000000 – 3000000 = 3000000micro joule = 3J..Thanks and RegardsAdarsh